OpenStudy (anonymous):

Solve. log[16]^(8)=x Is 0 correct?

OpenStudy (unklerhaukus):

\[\log_{16}8=x\]?

OpenStudy (anonymous):

Yes

OpenStudy (unklerhaukus):

do you know the change of base formula? \[\log_ba=\frac{\log_n a}{\log_nb}\]

OpenStudy (unklerhaukus):

change the base, to base 2

OpenStudy (anonymous):

thank you

OpenStudy (unklerhaukus):

and use \[8=2^3\qquad\qquad16=2^4\]